OpenStudy (anonymous):

can sum1 really help me n not leave? Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

@Agent_Sniffles

OpenStudy (anonymous):

slope =(y2-y1)/(x2-x1) -3 = (5-y)/(7-6)

OpenStudy (anonymous):

Now, solve for y

OpenStudy (anonymous):

how do u get the y?

OpenStudy (anonymous):

@s

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

(6, y) and (7, 5)

OpenStudy (anonymous):

7-6=1, so solve -3 = 5-y

OpenStudy (anonymous):

wats y?

OpenStudy (anonymous):

5-(-7) -2?

OpenStudy (anonymous):

y1--->y

OpenStudy (anonymous):

5-y?

OpenStudy (anonymous):

slope =(y2-y1)/(x2-x1) -3 = (5-y)/(7-6) -3=5-y y=8

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can u help with more?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

n Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of two thirds

OpenStudy (anonymous):

@sauravshakya